Lance Armstrong can ride 162 miles on flat ground in 6 hours with a good breeze at his back. it takes him 10 hours to go 90 miles with the same breeze working against him. How fast is lance going on a bike and how fast is the wind speed?
Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

There's no simple relationship between bike speed and windspeed.

Answer by mananth(16946)

bike speed =x mph
wind speed given speed =y mph
against wind speed given 10 hours
with wind speed given 6 hours

Distance against 90 miles distance with 162 miles
t=d/r against wind speed given -
90.00 / ( x - y )= 10.00
10.00 ( x - y ) = 6.00
10.00 x - 10.00 y = 90.00 ....................1

162.00 / ( x + y )= 6.00
6.00 ( x + y ) = 162.00
6.00 x + 6.00 y = 162.00 ...............2
Multiply (1) by 3.00
Multiply (2) by 5.00
we get
30.00 x + -30.00 y = 270.00
30.00 x + 30.00 y = 810.00
60.00 x = 1080.00
/ 60.00
x = 18.00 mph

plug value of x in (1) y
10.00 x -10.00 y = 90.00
180.00 -10.00 -180.00 = 90.00
-10.00 y = 90.00
-10.00 y = -90.00 mph
y = 9.00
bike speed 18.00 mph
wind speed given speed 9.00 mph

m.ananth@hotmail.ca